OpenStudy (anonymous):

-2(5x-3(2x+(x+y))+3)

OpenStudy (anonymous):

8x+6y-6

OpenStudy (saifoo.khan):

Solve the Core brackets first.

OpenStudy (luffingsails):

Morning Allyza. The key to simplification on these types of problems is as Saifoo put it. Start with the core brackets and then work out. In this case, I would start with: (2x+(x+y) This simplifies to: 2x+x+y = 3x+y Now I'm left with -2(5x-3(3x+y)+3), so I am going to focus on 3(3x+y) -3(3x+y) = -9x-3y, which gives me: -2(5x-9x-3y+3) = -2(-4x-3y+3) (after I resolve like terms) Now, I can bring in the -2 and do my final simplification -2(-4x-3y+3) = 8x+6y+6 Hope that helps and makes sense.

OpenStudy (anonymous):

The answer is 8x+6y-6 -2(5x-3(2x+(x+y))+3) -2(5x-3(2x+ x+y)+3) -2(5x-3(3x+y)+3) -2(5x-9x -3y +3) -10x +18x +6y - 6 8x + 6y - 6 Always work from the core out. The core here is (x + y)